# Optimal parallelisation strategies for flat histogram Monte Carlo sampling

### v2

@@ Abstract
Flat histogram methods, such as Wang–Landau sampling, provide a means for high-throughput calculation of phase diagrams of atomistic/lattice model systems. Many parallelisation schemes with varying degrees of complexity have been proposed to accelerate such sampling simulations. In this study, several widely used schemes are benchmarked—both in isolation and in combination—to establish best practice. The schemes studied include energy domain decomposition with both static sizing of energy sub-domains, as well as a dynamic sub-domain sizing scheme which we propose. We also assess the benefits both of replica exchange and of including multiple random walkers per sub-domain, to determine which factors have the largest impact on parallel efficiency. Additionally, the influence of energy sub-domain overlap regions is discussed. As an illustrative test case, we implement and apply the aforementioned strategies to a lattice-based model describing the internal energies of the AlTiCrMo refractory high-entropy superalloy, which is understood to crystallographically order into a B2 (CsCl) structure with decreasing temperature. We find that—while all of the proposed strategies confer a non-negligible speedup—parallelisation across energy domains which are non-uniform in size offers the most appreciable performance improvements. This work offers concrete recommendations for which parallelisation strategies should be prioritised to optimally accelerate flat-histogram Monte Carlo simulations.
Keywords: Parallelism, Monte Carlo methods, Wang–Landau sampling, Alloys, Phase diagrams, Thermodynamics
@@ I Introduction
Building on this concept, the Wang–Landau (WL) [10, 11, 12, 13, 14] sampling algorithm can be used to obtain the density of states (DOS) of a given model in the canonical ensemble, which can then be used to compute physical properties such as specific heat and internal energy—both as functions of temperature—as a simple post-processing step. The WL algorithm samples configurations in inverse proportion to the DOS, an initial estimate of which is iteratively refined during the sampling process until a uniform sampling distribution in energy is obtained. In this way, from a single simulation, one obtains both a set of samples which span the entire energy domain of interest, and the DOS over that same energy domain. WL sampling has found utility across a range of research areas, including the study of phase transitions [15, 16, 17], polymer physics [18, 19, 20], solubility prediction [21, 22], magnetic materials [23, 24, 25, 26, 27, 28], and surface adsorption (adhesion of atoms/ions/molecules to surfaces) [29, 30, 31]. Methods to update the statistical temperature of the system instead of the DOS within WL have also been proposed and used to develop an algorithm equivalent of WL for molecular dynamics known as statistical temperature molecular dynamics (STMD) [32].
@@
This paper aims to identify the most efficient parallel implementation of the WL algorithm, with a particular focus on iterative runtime load balancing. The context of our work is the need to perform a great many high throughput WL calculation with varying Hamiltonian parameters and material compositions. We are hence primarily concerned with maximising per-core parallel efficiency rather than reaching the highest possible speedup or core count for a single WL calculation. A key development of the present work is the demonstration that, despite the challenging nature of achieving perfectly optimal load balancing, it can be closely approached by dynamically adjusting the size of each energy sub-domain after each WL iteration, thereby leading to improved computational performance over long simulation runs. We also demonstrate how the performance of the algorithm can dramatically depend on several factors, including: the choice of the number of energy sub-domains and their size; the number of walkers within each energy sub-domain; and the size of the overlap between neighbouring energy sub-domains.
The remainder of this manuscript is structured as follows. First, in Sec. II.1 we give an overview of the WL sampling algorithm. Subsequently, in Sec. II.2, we discuss the various means by which the algorithm may be parallelised. Following on from this, in Sec. II.3 we present the model system used for benchmarking—a high-entropy alloy—and how it was implemented, along with a brief reprise of metrics for assessing parallel efficiency with respect to number of sub-domains and walkers. Proceeding in Sec. III, we present results comparing the efficiency of parallelisation schemes in the context of a benchmark model system, both in isolation and in combination. Finally, in Sec. IV, we conclude with recommendations on which parallelisation scheme(s) to prioritise when implementing flat histogram MC methods. Throughout the paper we will note where we expect conclusions to be specific to the system and where they are more generally applicable.
@@ II.1 Wang–Landau (WL) Sampling
WL sampling [10, 1] is a flat energy histogram technique with broad applicability. For a given model system—with an associated Hamiltonian that defines the energy as a function of the model’s configuration—the algorithm determines the DOS, from which thermodynamic quantities can subsequently be calculated for any temperature. This is in contrast to the simpler Metropolis MC algorithm, in which a separate simulation is required at every discrete temperature of interest. In order to obtain the DOS, we must first define an accessible energy domain of interest. This is perhaps most easily found by performing a simulation using the Metropolis algorithm to determine the extrema of the energy at the highest and lowest temperatures of interest. Using the energies found as a guide, we can choose an energy domain which is uniformly divided into $N$ bins of equal size, with each bin representing an energy macrostate containing multiple model system configurations. Given a model in which microstates are discrete—such as the lattice-based model in this study—the partition function $Z$ is expressed as a sum over all microstates. This can be re-written as a sum over all the discrete energies $E$ [1], i.e.
@@
where $E_{j}$ is the energy of the centre of the $j^{\text{th}}$ energy bin, and $g(E_{j})$ is the DOS corresponding to the $j^{\text{th}}$ energy bin. (Note that the first summation appearing in Eq. 1 is the usual summation over microstates, and that the second summation of Eq. 1 is taken over all energies accessible to the system. This is in contrast to the approximate summation of Eq. 2, which is a sum over the newly-defined energy bins, taken to represent macrostates.)
@@
$P_{n\rightarrow m}=\begin{cases}\;\frac{g(E_{n})}{g(E_{m})},&\quad g(E_{n})<g(E_{m})\\ \;1,&\quad g(E_{n})\geq g(E_{m}),\end{cases}$ (3)
@@
Independently of any parallelisation strategy, there have been efforts to optimise both the algorithm’s ‘modification factor’ and ‘flatness criterion’ [36, 37, 38, 39]. The WL algorithm was quickly improved by the introduction of energy domain decomposition into sub-domains [10]. This decomposition allowed for independent random walks to be performed either in serial or concurrently. Further optimisation can be achieved by dynamically adjusting the size of each energy sub-domain, although achieving maximally optimal load balancing is challenging due to the complex nature of the free energy landscape for a given system, several methods have been proposed for the dynamic adjustment of sub-domain size within a WL simulation [40, 41, 42]. These methods typically rely on restricting the energy domain over which a single WL random walk occurs. Once the flatness criterion has been met on a subset of the total energy domain, that portion of the energy domain is excluded from the current available energy for sampling during the iteration. With the advent of more accessible multi-process computing, these methods can be further expanded by combining non-uniform energy sub-domains with adaptive energy sub-domain sizing, instead of just modifying the total sampling energy domain for a single WL instance.
@@ II.2 Parallelisation Strategies
Building on these ideas, all parallelisation schemes compared in this work make use of energy domain decomposition in some form. This involves subdividing the total energy domain considered in the simulation into smaller energy sub-domains, each of which is sampled by one or more independent WL instances, i.e. a random walker or set of walkers. In our implementation, each random walker constitutes an independent CPU process concurrently evolved throughout the simulation. A conceptual illustration of the two different energy domain decomposition schemes considered in this work is given in panels (a) and (b) of Figure 1.
@@
$P_{n\rightarrow m}=\begin{cases}\;1,&\quad(E_{m}-E_{T})^{2}-(E_{n}-E_{T})^{2}\leq 0\\ \;0,&\quad(E_{m}-E_{T})^{2}-(E_{n}-E_{T})^{2}>0,\end{cases}$ (8)
@@ II.2.2 Non-Uniform Energy Domain Decomposition
The time taken for each WL iteration is given by the time to converge of the slowest-converging sub-domain, since each walker must first converge their local (sub-domain) DOS histogram before the WL iteration can finish and create a combined global DOS. In general, the time taken to converge each sub-domain is non-uniform across the total energy domain, tending to take longer to converge around phase transitions or at low energy in the presence of degenerate microstates. In the former case, systems may exhibit entropic bottlenecks which are difficult to sample, and in the latter case the number of accessible configurations necessitates long sampling times. This non-uniform convergence time can be somewhat mitigated by altering the size of each WL energy sub-domain, however optimal load balancing is impossible a priori, as that requires the converged DOS states, which is the quantity being computed via simulation. By adjusting the individual size of each sub-domain, one can obtain an energy decomposition which is non-uniform in energy. This is done before the first iteration via a pre-sampling step, which evolves the WL algorithm until each energy bin across the entire energy domain has been visited by at least one walker per sub-domain a thousand times. Performing a pre-sampling step that is much longer than this is inefficient, since the simulation is progressing with sub-domain sizes which are non-optimal. This pre-sampling step is performed with uniformly sized sub-domains and lasts on the order of one WL iteration. The time taken per sub-domain is used to obtain the initial non-uniform energy domain decomposition. The details of the explicit means by which this initial non-uniform energy domain decomposition is obtained are discussed below, in Sec. II.2.4. A conceptual illustration of a non-uniform energy domain decomposition scheme is given in Figure 1, panel (b).
@@ II.2.4 Dynamic Load Balancing
where $d_{i,k}$ is the size of sub-domain $i$ after iteration $k$ as a fraction of the total number of bins, and $a_{k}$ and $b_{k}$ are scaling factors. To ensure the entire energy domain is covered, the $\sum_{i}d_{i,k+1}=N$. The purpose of $a_{k}$ and $b_{k}$ is to introduce memory (i.e. incremental changes to previous sub-domain sizing) into the size of the energy sub-domains in order to reduce overcompensating after an outlier in convergence time. With each subsequent WL iteration $k$, $b_{k}$ decreases such that the sub-domain adjustment decreases with each iteration. Explicitly, in this work, $a_{k}$ and $b_{k}$ were chosen according to
$a_{k}$ $=1-b_{k},$ (11)
$b_{k}$ $=\alpha\cdot\beta^{k},$
$a_{0}$ $=1-b_{0},$
$b_{0}$ $=\alpha\cdot\beta^{0},$
@@
where $\alpha$ controls the memory, and $\beta$ controls the adjustment rate of sub-domain size. We find $\alpha=0.2$ and $\beta=0.9$ to be a reliable choice of parameters for use in the context of the present work. Of note, the initial update following the pre-sampling step is performed using a uniformly initialised $d$ with $b=1$. Additionally, in this work, to prevent zero-bin energy sub-domains, a minimum sub-domain size is imposed of 2% of the total number of bins across the entire energy domain.
@@
After such a sub-domain adjustment, walkers may no longer be in their designated sub-domain and as such they are moved into their appropriate energy domain, as outlined at the start of this section, i.e. by performing (and restarting as necessary) an MC simulation with acceptance probability as defined in Eq. 8. (Note that accumulation of the histogram(s) and refinement of the DOS are paused during this adjustment process.) The sizes of energy overlap regions are also re-computed for the adjusted sub-domain sizing. We note here that a related approach has previously been explored, in which a single load balancing step was performed following the pre-sampling step [33], but where there was no further adjustment of energy sub-domain sizing following this initial load-balancing.

### v3

@@ Abstract
Flat histogram methods, such as Wang–Landau sampling, provide a means for high-throughput calculation of phase diagrams of atomistic/lattice model systems. Many parallelisation schemes with varying degrees of complexity have been proposed to accelerate such sampling simulations. In this study, several widely used schemes are benchmarked—both in isolation and in combination—to establish best practice. The schemes studied include energy domain decomposition with both static sizing of energy sub-domains, as well as a dynamic sub-domain sizing scheme which we propose. We also assess the benefits both of replica exchange and of including multiple random walkers per sub-domain, to determine which factors have the largest impact on parallel efficiency. Additionally, the influence of energy sub-domain overlap regions is discussed. As illustrative test cases, we implement and apply the aforementioned strategies to a lattice-based model describing the internal energy of a substitutional alloy, studying the AlTiCrMo refractory high-entropy superalloy as well as the binary CuZn system, both of which crystallographically order into a B2 (CsCl) structure with decreasing temperature. We find that—while all of the proposed strategies confer a non-negligible speedup—parallelisation across energy domains which are non-uniform in size offers the most appreciable performance improvements. This work offers concrete recommendations for which parallelisation strategies should be prioritised to optimally accelerate flat-histogram Monte Carlo simulations.
Keywords: Parallelism, Monte Carlo, Sampling, Alloys, Thermodynamics
@@ I Introduction
Building on this concept, the Wang–Landau (WL) [10, 11, 12, 13, 14] sampling algorithm can be used to obtain the density of states (DOS) of a given model in the canonical ensemble, which can then be used to compute physical properties such as specific heat and internal energy—both as functions of temperature—as a simple post-processing step. The WL algorithm samples configurations in inverse proportion to the DOS, an initial estimate of which is iteratively refined during the sampling process until a uniform sampling distribution in energy is obtained. In this way, from a single simulation, one obtains both a set of samples which span the entire energy domain of interest, and the DOS over that same energy domain. WL sampling has found utility across a range of research areas, including the study of phase transitions [15, 16, 17, 18], polymer physics [19, 20, 21], solubility prediction [22, 23], magnetic materials [24, 25, 26, 27, 28, 29], and surface adsorption (adhesion of atoms/ions/molecules to surfaces) [30, 31, 32]. Methods to update the statistical temperature of the system instead of the DOS within WL have also been proposed and used to develop an algorithm equivalent of WL for molecular dynamics known as statistical temperature molecular dynamics (STMD) [33].
@@
This paper aims to identify the most efficient parallel implementation of the WL algorithm, with a particular focus on iterative runtime load balancing. The context of our work is the need to perform a great many high-throughput WL calculations with varying Hamiltonian parameters and material compositions. We are hence primarily concerned with maximising per-core parallel efficiency rather than reaching the highest possible speedup or core count for a single WL calculation. A key development of the present work is the demonstration that, despite the challenging nature of achieving perfectly optimal load balancing, it can be closely approached by dynamically adjusting the size of each energy sub-domain after each WL iteration, thereby leading to improved computational performance over long simulation runs. We also demonstrate how the performance of the algorithm can dramatically depend on several factors, including: the choice of the number of energy sub-domains and their size; the number of walkers within each energy sub-domain; and the size of the overlap between neighbouring energy sub-domains.
The remainder of this manuscript is structured as follows. First, in Sec. II.1 we give an overview of the WL sampling algorithm. Subsequently, in Sec. II.2, we discuss the various means by which the algorithm may be parallelised. Following on from this, in Sec. II.3 we present the two model systems used for benchmarking—a high-entropy alloy and a binary alloy—and how they were implemented, along with a brief reprise of metrics for assessing parallel efficiency with respect to number of sub-domains and walkers. Proceeding in Sec. III, we present results comparing the efficiency of parallelisation schemes in the context of a benchmark model system, both in isolation and in combination. Finally, in Sec. IV, we conclude with recommendations on which parallelisation scheme(s) to prioritise when implementing flat histogram MC methods. Throughout the paper we will note where we expect conclusions to be specific to the system and where they are more generally applicable.
@@ II.1 Wang–Landau (WL) Sampling
WL sampling [10, 1] is a flat energy histogram technique with broad applicability. For a given model system—with an associated Hamiltonian that defines the energy as a function of the model’s configuration—the algorithm determines the DOS, from which thermodynamic quantities can subsequently be calculated for any temperature. This is in contrast to the simpler Metropolis MC algorithm, in which a separate simulation is required at every discrete temperature of interest. In order to obtain the DOS, we must first define an accessible energy domain of interest. This is perhaps most easily found by performing a simulation using the Metropolis algorithm to determine the extrema of the energy at the highest ($T_{\text{max}}$) and lowest ($T_{\text{min}}$) temperatures of interest. Note that the lowest temperature of interest is not necessarily one at which the ground state of the model system has non-vanishing statistical weight if, for example, the simulation is being performed to locate phase transitions at higher temperatures. Using the energies found as a guide, we can choose an energy domain which is uniformly divided into $N$ bins of equal size, with each bin representing an energy macrostate containing multiple model system configurations. Given a model in which microstates are discrete—such as the lattice-based models in this study—the partition function $Z$ is expressed as a sum over all microstates. This can be re-written as a sum over all the discrete energies $E$ [1], i.e.
@@
where $E_{j}$ is the energy of the centre of the $j^{\text{th}}$ energy bin, and $g(E_{j})$ is the DOS corresponding to the $j^{\text{th}}$ energy bin. (Note that the first summation appearing in Eq. 1 is the usual summation over microstates, and that the second summation of Eq. 1 is taken over all energies accessible to the system. This is in contrast to the approximate summation of Eq. 2, which is a sum over the newly-defined energy bins, taken to represent macrostates.) If the system being studied is particularly discrete, then it is possible for energy states accessible to the system to be spaced further apart than the width of the WL energy bins. This can result in energy bins that have no energy states and are therefore not possible to visit, which would interfere with WL sampling. To remedy this, one can simply adjust the binning such that there are no bins without available energy states. In general, we expect this to not be an issue for systems that will benefit from the parallelisation schemes outlined within this study.
@@
$P_{n\rightarrow m}=\cases{\;}\frac{g(E_{n})}{g(E_{m})},&\quad g(E_{n})<g(E_{m})\\ \;1,&\quad g(E_{n})\geq g(E_{m}),$ (3)
@@
Independently of any parallelisation strategy, there have been efforts to optimise both the algorithm’s ‘modification factor’ and ‘flatness criterion’ [37, 38, 39, 40]. The WL algorithm was quickly improved by the introduction of energy domain decomposition into sub-domains [10]. This decomposition allowed for independent random walks to be performed either in serial or concurrently. Further optimisation can be achieved by dynamically adjusting the size of each energy sub-domain; however, achieving maximally optimal load balancing is challenging due to the complex nature of the free energy landscape for a given system, and several methods have been proposed for the dynamic adjustment of sub-domain size within a WL simulation [41, 42, 43]. These methods typically rely on restricting the energy domain over which a single WL random walk occurs. Once the flatness criterion has been met on a subset of the total energy domain, that portion of the energy domain is excluded from the current available energy for sampling during the iteration. With the advent of more accessible multi-process computing, these methods can be further expanded by combining non-uniform energy sub-domains with adaptive energy sub-domain sizing, instead of just modifying the total sampling energy domain for a single WL instance. It is worth mentioning that the systems that will benefit the most from the schemes to be outlined in the subsequent section are those that are difficult to converge. If the model system being studied rapidly converges to a desired tolerance, the benefit of parallelisation can be significantly diminished.
Of interest is also the application of GPUs for the acceleration of WL sampling, where thousands of concurrent walkers can be run on a single GPU [44]. Further improvements to WL sampling on CPUs were made by the development of scalable replica exchange implementations [45], with our implementation of replica exchange being discussed in the subsequent section. Another source of potential acceleration stems from the fact that $\log(g(E))$ typically follows a parabola centred on its maximum for alloys that exhibit one phase transition [46]. However, for an arbitrary system the form of $\log(g(E))$ is not known. If one knows that the underlying $\log(g(E))$ is parabolic then in theory this quantity could be initialised to a parabola or one could try to dynamically fit a parabola based upon the DOS obtained during pre-sampling. While this has potential for WL sampling, it would not apply to alloys that exhibit more than one phase transition and as such is not considered in this work.
@@ II.2 Parallelisation Strategies
Building on these ideas, all parallelisation schemes compared in this work make use of energy domain decomposition in some form. This involves subdividing the total energy domain considered in the simulation into smaller energy sub-domains, each of which is sampled by one or more independent WL instances, i.e. a random walker or set of walkers. In our implementation, each random walker constitutes an independent CPU process concurrently evolved throughout the simulation. A conceptual illustration of the two different energy domain decomposition schemes considered in this work is given in panels (a) and (b) of Figure 1. It should also be stressed that implementation complexity is a key factor to consider when selecting parallelisation scheme(s). Therefore, below, once the parallelisation schemes have been outlined, we shall discuss this key aspect.
@@
$P_{n\rightarrow m}=\cases{\;}1,&\quad(E_{m}-E_{T})^{2}-(E_{n}-E_{T})^{2}\leq 0\\ \;0,&\quad(E_{m}-E_{T})^{2}-(E_{n}-E_{T})^{2}>0,$ (8)
@@ II.2.2 Non-Uniform Energy Domain Decomposition
The time taken for each WL iteration is given by the time to converge of the slowest-converging sub-domain, since each walker must first converge their local (sub-domain) DOS histogram before the WL iteration can finish and create a combined global DOS. In general, the time taken to converge each sub-domain is non-uniform across the total energy domain, tending to take longer to converge around phase transitions or at low energy in the presence of degenerate microstates. In the former case, systems may exhibit entropic bottlenecks which are difficult to sample, and in the latter case the number of accessible configurations necessitates long sampling times. This non-uniform convergence time can be somewhat mitigated by altering the size of each WL energy sub-domain, however optimal load balancing is impossible a priori, as that requires the converged DOS states, which is the quantity being computed via simulation. By adjusting the individual size of each sub-domain, one can obtain an energy decomposition which is non-uniform in energy. This is done before the first iteration via a pre-sampling step, which evolves the WL algorithm until each energy bin across the entire energy domain has been visited by at least one walker per sub-domain a thousand times. We selected this number as a compromise between excessive simulation time being spent in the pre-sampling step and ensuring that every bin has sufficient visits such that the initial DOS estimate is reasonable. We find that the selection of this parameter is not crucial, since an estimate of the DOS must be built up from an initial state of $\log(g(E))$ which takes an appreciable portion of simulation time regardless, and so the parameter is chosen to err on the side of caution to ensure that all bins are appropriately sampled. Performing a pre-sampling step that is much longer than this is inefficient, since the simulation is progressing with sub-domain sizes which are non-optimal. This pre-sampling step is performed with uniformly sized sub-domains and lasts on the order of one WL iteration. The time taken per sub-domain is used to obtain the initial non-uniform energy domain decomposition. The details of the explicit means by which this initial non-uniform energy domain decomposition is obtained are discussed below, in Sec. II.2.4. A conceptual illustration of a non-uniform energy domain decomposition scheme is given in Figure 1, panel (b).
@@ II.2.4 Dynamic Load Balancing
where $d_{i,k}$ is the size of sub-domain $i$ after iteration $k$ as a fraction of the total number of bins, and $a_{k}$ and $b_{k}$ are scaling factors. To ensure the entire energy domain is covered, $\sum_{i}d_{i,k+1}=N$. The purpose of $a_{k}$ and $b_{k}$ is to introduce memory (i.e. incremental changes to previous sub-domain sizing) into the size of the energy sub-domains in order to reduce overcompensating after an outlier in convergence time. With each subsequent WL iteration $k$, $b_{k}$ decreases such that the sub-domain adjustment decreases with each iteration. Explicitly, in this work, $a_{k}$ and $b_{k}$ were chosen according to
$aligned a_{k}&=1-b_{k},\\ b_{k}&=\alpha\cdot\beta^{k},\\ a_{0}&=1-b_{0},\\ b_{0}&=\alpha\cdot\beta^{0},$ (11)
@@
where $\alpha$ controls the memory, and $\beta$ controls the adjustment rate of sub-domain size. We find $\alpha=0.2$ and $\beta=0.9$ to be a reliable choice of parameters for use in the context of the present work. While results are sensitive to the choice of these parameters, their only effect is the rate at which the optimal domain sizes are found. In general, one should take care to choose $\alpha$ and $\beta$ such that the rate of change of domain size decreases appreciably to prevent excessive oscillation of domain size. Of note, the initial update following the pre-sampling step is performed using a uniformly initialised $d$ with $b=1$. Additionally, in this work, to prevent zero-bin energy sub-domains, a minimum sub-domain size is imposed of 2% of the total number of bins across the entire energy domain.
@@
After such a sub-domain adjustment, walkers may no longer be in their designated sub-domain and as such they are moved into their appropriate energy domain, as outlined at the start of this section, i.e. by performing (and restarting as necessary) an MC simulation with acceptance probability as defined in Eq. 8. (Note that accumulation of the histogram(s) and refinement of the DOS are paused during this adjustment process.) The sizes of energy overlap regions are also re-computed for the adjusted sub-domain sizing. We note here that a related approach has previously been explored, in which a single load balancing step was performed following the pre-sampling step [34], but where there was no further adjustment of energy sub-domain sizing following this initial load-balancing. We stress however, that this per-iteration adjustment has negligible computational overhead as compared to the total runtime of the simulation ($0.1\%$, taking $\sim$15 ms per call in our implementation).
Figure 2: Plots of real-space effective pair interactions, $V_{\alpha\alpha^{\prime}}^{(n)}$, as a function of coordination shell number, $n$, recovered from the reciprocal space $S^{(2)}_{\alpha\alpha^{\prime}}({\bf k})$ data for (a) AlTiCrMo and (b) CuZn. The notation $V_{\alpha\alpha^{\prime}}^{(n)}$ indicates the interaction between chemical species $\alpha$ and $\alpha^{\prime}$ on coordination shell $n$, i.e. at $n$th nearest neighbour distance. For both alloys, it can be seen that the comparative strength of interactions tails off quickly with increasing distance and is particularly rapid for CuZn. Note that, for the CuZn binary, the Cu-Cu and Zn-Zn effective pair interactions are equal on all coordination shells due to the choice of gauge [52].
II.2.5 Implementation Complexity
We now comment on the implementation complexity of each of the outlined parallelisation schemes and features. We highlight that, provided all parallelism features are implemented one after another, they form natural extensions of each other, simplifying the overall implementation. In order, the complexity of implementation of the various features are summarised as follows:
  • Energy domain decomposition (uniform). Implementation of uniform domain decomposition is straightforward as each sub-domain can be treated as an independent WL instance with its own histograms, and therefore requiring no changes to the core WL algorithm. The two main implementation requirements are: ensuring all WL instances share the same modification factor and flatness criterion, and enforcing that each WL instance remains within its assigned energy interval.
  • Global DOS construction. Constructing a global DOS is the primary non-trivial component of uniform energy domain decomposition. A robust combination procedure is needed that is capable of identifying a joining point within the overlap region between neighbouring sub-domains (here based on $\beta(E)$), and applying a consistent offset between $\log g(E)$ segments so that they match at the selected joining point. If this is performed without sufficient care, artefacts in the resulting DOS can occur at the boundaries between sub-domains.
  • Energy domain decomposition (non-uniform). Non-uniform energy domain decomposition is more complex than the uniform case but retains the same overall structure. Each sub-domain remains as an independent WL instance and, as such, most uniform decomposition functionality can be reused with the only additional requirement being a method for measuring convergence time per sub-domain. These timings can then be used during a pre-sampling step to define new, non-uniform sub-domains, after which the existing functionality for keeping instances within their target domains can be reused.
  • Replica exchange. Replica exchange has moderate implementation complexity. It introduces explicit communication between neighbouring sub-domains via configuration swaps, requiring a synchronised swap protocol (proposal schedule, overlap energy eligibility checks, and accept/reject decisions). Care is needed to avoid communication deadlocks and to ensure exchanges occur only between eligible WL instances. Unlike overlap handling or DOS construction, replica exchange must be integrated directly into the sampling loop rather than added as an intermediate step between WL iterations.
  • Dynamic load balancing. While dynamic load balancing appears complex, it becomes a straightforward extension of non-uniform domain decomposition once this feature is implemented. The additional functionality required is a per-iteration mechanism to update sub-domain boundaries, and a method to map each updated sub-domain onto the corresponding region of the global DOS.
For the purposes of the present work, we have implemented these schemes in the open-source BraWl package [53].